OpenStudy (anonymous):

Determine the standard form of the equation of the line that passes through (6,-9) and (-8,6)

OpenStudy (anonymous):

first find the slope. and then use the equation \[\left( y - y _{1} \right) = m \left( x - x _{1} \right)\]

OpenStudy (anonymous):

\[m = \frac{ y _{2} - y _{1}}{ x _{2} - x _{1} }\]

OpenStudy (anonymous):

y+9=15/-14(x-6)